A machine costing $210,000 with a four-year life and an estimated $18,000 salvage value is installed in Luther Company’s factory on January 1. The factory manager estimates the machine will produce 480,000 units of product during its life. It actually produces the following units: 123,000 in 1st year, 123,500 in 2nd year, 119,900 in 3rd year, 123,600 in 4th year. The total number of units produced by the end of year 4 exceeds the original estimate—this difference was not predicted. (The machine must not be depreciated below its estimated salvage value.)

Compute depreciation for each year (and total depreciation of all years combined) for the machine under Straight-line depreciation.

Straight-Line Depreciation

Year

Depreciation Expense

1

2

3

4

Total

$0

Units of Production

Year

Depreciable Units

Depreciation per unit

Depreciation Expense

1

2

3

4

Total

$0

Compute depreciation for each year (and total depreciation of all years combined) for the machine under Double-declining-balance.

DDB Depreciation for the Period

End of Period

Year

Beginning of Period Book Value

Depreciation Rate

Depreciation Expense

Accumulated Depreciation

Book Value

1

%

$0

2

%

0

3

%

0

4

%

0

$0
Answer #1

Straight line Depreciation per year = (210000-18000)/4= 48,000

Year 1 48,000

Year 2 48,000

Year 3 48,000

Year 4 48,000

Units of production :

Depreciable units Depreciation per unit Depreciation expense
120000 0.4 48000
120000 0.4 48000
120000 0.4 48000
120000 0.4 48000

Double declining balance method

Beginning value Rate Depreciation Net book value Accumulated Depreciation
              2,10,000 50 105000 105000                                       1,05,000
              1,05,000 50 52500 52500                                       1,57,500
                 52,500 50 26250 26250                                       1,83,750
                 26,250 50 8250 18000                                       1,92,000

Rate if depreciation computed as = 100/4*2= 50